OpenStudy (joshoyen):

HELP Solve by factoring. 5x^2+47x+56=0

OpenStudy (xapproachesinfinity):

use the ac method 5x56=280 find the factor of 280 that sum up to 47

OpenStudy (joshoyen):

hmm, im still confused :L

OpenStudy (joshoyen):

oh wait is it 7 x 40

OpenStudy (xapproachesinfinity):

will any second degree poly can be written like a(x+b)(x+c) the goal here is to a,b,c

OpenStudy (xapproachesinfinity):

but ac method don't worry about a just b and c

OpenStudy (xapproachesinfinity):

yes that is correct

OpenStudy (xapproachesinfinity):

40x7=280 and 40+7=47 good

OpenStudy (xapproachesinfinity):

now lets rewrite the equation \[5x^2+47x+56=5x^2+40x+7x+56=(5x^2+40)+(7x+56)\]

OpenStudy (xapproachesinfinity):

now what you can see is 5x^2and 40x have 5x in common

OpenStudy (xapproachesinfinity):

error in parenthesis i should have (5x^2+40x) not (5x^2+40) forget x

OpenStudy (xapproachesinfinity):

and we can also see that 7x and 56 have 7 in common (common factor) so we can now do this \[=5x(x+8)+7(x+8)=(5x+7)(x+8)\]

OpenStudy (xapproachesinfinity):

so we got (5x+7)(x+8)

OpenStudy (joshoyen):

ohhhh , okay now i get it

OpenStudy (joshoyen):

alright thankyou!

OpenStudy (xapproachesinfinity):

you welcome

OpenStudy (joshoyen):

wait can we do one more equation?

OpenStudy (joshoyen):

7a^2 - a - 8 = 0